# Expected Information Gain Estimation via Density Approximations: Sample Allocation and Dimension Reduction

### v1

@@ 4.1 Sample allocations and convergence rates
To investigate the effect of sample allocation, we set the ratio between the evaluation samples and training samples to be $L^{p}$, with $p=\frac{1}{8},\frac{1}{3},\frac{3}{4}$ respectively, and each experiment is repeated 100 times. In the first study, we assume the model is known and the likelihood function can be evaluated exactly, and we estimate only the marginal distribution of $Y$. That is, we use $\widehat{\mathrm{EIG}}_{\mathrm{m}}$. We also compare the performance of TM-based estimators with that of the NMC estimator, where the NMC estimator is computed using the optimal sample allocation discussed in Feng and Marzouk [23]. We then plot the estimated EIG along with their bias, variance, and mean squared error (MSE) for the different estimators in Figure 1 and 2. We observe that when we set the ratio of $M$ and $N$ to be proportional to $L^{1/3}$, the MSE decreases most rapidly as the total number of samples increases. The empirical convergence rate also aligns with the theory presented in the previous section. Next, we show in Figures 3 and  4 that by fixing the prior and using TM to approximate the posterior density, i.e., setting the EIG estimator to $\widehat{\mathrm{EIG}}_{\mathrm{pos}}$, we achieve the same rate.
@@
We then study the likelihood-free case, where we estimate either the likelihood and the marginal densities and obtain $\widehat{\mathrm{EIG}}_{\mathrm{lik}}$ (Figure 5), or the posterior and prior densities and obtain $\widehat{\mathrm{EIG}}_{\mathrm{pr}}$ (Figure 7). As shown in the Figure 6 and 8, in both cases, the convergence rate of the MSE for the estimator matches the theoretical rate in the asymptotic regime.
@@
We then present a few observations: The convergence rate of the MSE with sample allocation $\frac{M}{N}\sim O(L^{1/3})$ is of order $L^{-1}$ in the asymptotic region across all four cases. For $L^{p}$, when $p>\frac{1}{2}$, we observe that the convergence rate of the MSE is considerably slower than for $p\leq\frac{1}{2}$ in all four cases, which supports our theoretical results developed in Section  2.3. From the violin plot, we see that all the TM-based results have smaller variance than NMC, and with allocation ratios of $L^{1/8}$ and $L^{1/3}$, TM produces good results even with small sample sizes. Another noteworthy observation is that $\widehat{\mathrm{EIG}}_{\mathrm{lik}}$ and $\widehat{\mathrm{EIG}}_{\mathrm{pr}}$ yield exactly the same values (see Figures 5 and  7). This occurs because we are in the linear Gaussian case, and if we assume the mean has been shifted to 0, then the map can be computed from covariance matrix of the joint samples. In other words, we can obtain $\widehat{\pi}_{Y}$, $\widehat{\pi}_{Y|X}$, $\widehat{\pi}_{X}$, and $\widehat{\pi}_{X|Y}$ directly from the sample covariance matrix of the joint distribution. Therefore, we have exactly $\frac{\widehat{\pi}_{X|Y}(x|y)}{\widehat{\pi}_{X}(x)}=\frac{\widehat{\pi}_{Y|X}(y|x)}{\widehat{\pi}_{Y}(y)}$ for any $x$ and $y$.
@@ 4.2 Transport-based estimators of focused and full EIG
We then study a nonlinear example, Mössbauer spectroscopy, to demonstrate the flexibility of transport maps. Mössbauer spectroscopy is a technique that leverages the Mössbauer effect, which refers to the nearly recoil-free emission and absorption of gamma radiation in solids. This technique is used to discover isomer shifts, quadrupole splitting, and magnetic Zeeman splitting. A detailed description of the model can be found in Feng and Marzouk [23] and Mössbauer [44]. The parameters in this model are “center”, “width”, “height”, and “offset”. The absorption peak is modeled using a Lorentzian profile, as per the standard parameterization described in [28]. This allows the number of detector counts $y_{i}$ at a given velocity $d$ to be expressed by the following equation,
@@
All densities are estimated using $N$ pairs of samples. We set the “ground truth” to be the EIG estimator obtained using NMC. To obtain a consistent estimator of EIG using Monte Carlo for the focused case, we employed layered multiple importance sampling (LMIS) from Feng and Marzouk [23] with $1.58\times 10^{6}$ samples (5623 for the outer loops and 281 for the inner loops). For the full case, we used the NMC algorithm with $10^{10}$ samples, where the allocation of samples between the inner and outer loops is determined using the optimal allocation proposed in Feng and Marzouk [23]. In this case, we have $3.83\times 10^{6}$ samples for the inner loop and 2610 for the outer loop. We repeated the procedure 10 times to calculate the mean and variance of the EIG estimators obtained using LMIS and NMC. The mean of the Monte Carlo estimators in the focused and full cases is $1.54$ and $4.52$ respectively, with corresponding variances of $4.73\times 10^{-5}$ and $4.59\times 10^{-4}$. We then obtained EIG estimators using transport maps with an increasing total number of samples: 500, 5000, and 50000, and with the number of training and evaluation samples split according to the optimal sample allocation $M/N\sim O(L^{1/3})$, as derived in the previous section. We repeated the experiments, and obtained $30$ transport map-based EIG estimators. We present the numerical results of both cases in Figures 10 and 11.
@@
Finally, we present the joint distribution after applying the transport maps in Figure 12(a). As seen, the joint distribution after the push-forward map is very close to a standard normal distribution. To further illustrate the Gaussianity of the push-forward samples, we present the quantile-quantile plot (Q–Q plot) of the push-forward samples against that of a standard normal. To be precise, we generate 20 random vectors of unit length, and project 7-dimensional data onto 20 random directions, respectively. We then plot the quantile of these random projections against that of a standard normal distribution. If the push-forward distribution is exactly Gaussian, the Q–Q plot of the random projections would align with the black line (see Box 2.4 in Santambrogio [57]). As shown in Figure 12(b), only the tails of the push-forward samples deviate slightly, demonstrating the effectiveness of transport maps in transforming highly non-Gaussian samples into a standard Gaussian distribution.
@@
Figure 11: Convergence of different EIG estimators in the full case: (a) $\widehat{\mathrm{EIG}}_{\mathrm{lik}}$; (b) $\widehat{\mathrm{EIG}}_{\mathrm{pr}}$; (c) $\widehat{\mathrm{EIG}}_{\mathrm{m}}$ and $\widehat{\mathrm{EIG}}_{\mathrm{pos}}$. The dashed line represents the “ground truth” obtained using NMC.
@@ 4.3 EIG estimation in high dimensions: linear elasticity inverse problem
In this problem, we compare the proposed dimension reduction method for EIG estimation, referred to as Conditional Mutual Information (CMI), to using alternative decompositions based on Principal Component Analysis (PCA) [29] and Canonical Correlation Analysis (CCA) [30]. We consider the problem of estimating mutual information between the Young’s modulus field in a wrench-shaped domain and a noisy observation of its displacement on the upper boundary after applying an external force to a subset of the boundary $\partial\Omega$. Following the setup in [2, 38, 58], the displacement $u:\Omega\rightarrow\mathbb{R}^{2}$ is related to the Young’s modulus $E:\Omega\rightarrow\mathbb{R}_{\geq 0}$ through the partial differential equation
@@
$\Sigma_{X}u^{\operatorname{PCA}}$ $=\lambda\Sigma_{X}u^{\operatorname{PCA}}$
$\Sigma_{Y}v^{\operatorname{PCA}}$ $=\lambda\Sigma_{Y}v^{\operatorname{PCA}}.$
@@
We then transform the samples $x^{i}$ and $y^{i}$ using these eigenvectors $\{u^{\operatorname{CMI}},v^{\operatorname{CMI}}\},\{u^{\operatorname{PCA}},v^{\operatorname{PCA}}\}$, and $\{u^{\operatorname{CCA}},v^{\operatorname{CCA}}\}$, respectively. The maps are trained using $N=2500$ samples. We then compute the Monte Carlo average using $M=10000$ evaluation samples to obtain the estimator $\frac{1}{M}\sum_{i=1}^{M}\frac{\widehat{\pi}_{X_{r}|Y_{s}}(x_{r}^{i}|y_{s}^{i})}{\pi_{X_{r}}(x_{r}^{i})}$. To illustrate the effectiveness of our proposed nonlinear dimension reduction method, we compute EIG as if the data were Gaussian. To be precise, we compute the EIG using the formula $\frac{1}{2}\log\frac{\det(\widehat{\Sigma}_{X})\det(\widehat{\Sigma}_{Y})}{\det(\widehat{\Sigma}_{X,Y})}$. Here $\widehat{\Sigma}_{X},\widehat{\Sigma}_{Y}$ and $\widehat{\Sigma}_{X,Y}$ denotes the sample covariance of $X,Y$, and the joint distribution of $X$ and $Y$, respectively. In practice, the covariance matrices are approximated using their empirical counterparts estimated using $10000$ samples. For reproducibility, the same experiment is then repeated $10$ times, and we show the error bars, representing two standard error of the mean in Figure 14.
@@
Figure 14: EIG estimates using TM and Gaussian, with truncated dimensions using CMI, PCA, and CCA for $r=s=1$ to $8$
@@ 5 Conclusion
In this paper, we study a key problem in Bayesian optimal experimental design — EIG estimation. We introduced the transport-map based method for EIG estimation in Section 2. This enables us to estimate EIG with both explicit and implicit models. In Section 2.3, we study the optimal sample allocation between the training and the evaluation samples for the purpose of minimizing the MSE. This is further verified with the linear Gaussian example presented in Section 4.1. In Section 3, we adopt the optimal dimension reduction technique from Baptista et al. [2], enabling EIG to be estimated in high-dimensional settings. The numerical example presented in Section 4 show that our proposed method performs similarly comparing to nested Monte Carlo in terms of the EIG but with a much cheaper cost. We also demonstrate the performance of our method in the high-dimensional setting in Section 4.3. The dimension reduction technique proposed in this paper outperform PCA, CCA, and its Gaussian counterpart in terms of EIG estimation.
@@ Appendix B Proofs of the main results
This quantity can be estimated using samples:
$\mathrm{EIG}\approx\frac{1}{M}\sum_{i=1}^{M}\log\frac{\pi_{Y|X}(y^{i}|x^{i})}{\pi_{Y}(y^{i})}.$
@@
We then study the term-wise covariance:
$\mathbb{C}ov\left(\frac{1}{M}\sum_{i=1}^{M}\log\pi_{Y|X}(y^{i}|x^{i}),\frac{1}{M}\sum_{i=1}^{M}\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right)$
$=$ $\frac{1}{M^{2}}\sum_{i=1,j=1}^{M}\mathbb{C}ov\left(\log\pi_{Y|X}(y^{i}|x^{i}),\log q_{Y}(y^{j};\widehat{\alpha}_{N})\right).$
Note that $\mathbb{C}ov\left(\log\pi_{Y|X}(y^{i}|x^{i}),\log q_{Y}(y^{j};\widehat{\alpha}_{N})\right)=0$ for $i\neq j$. Therefore the expression above can be simplified to
$\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{C}ov\left(\log\pi_{Y|X}(y^{i}|x^{i}),\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right)\sim O\left(\frac{1}{M}\right),$
as $\mathbb{C}ov\left(\log\pi_{Y|X}(y^{i}|x^{i}),\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right)\sim O(1)$. To conclude, we have
@@
Proof B.3 (Proof of Corollary 2.3).
@@ Appendix C Chi-square distribution
Suppose $X\in\mathbb{R}^{n}$ is normally distributed, i.e., $X\sim\mathcal{N}(\mu,\Sigma_{1})$, then the random variable $Y=(X-\mu)^{\top}\Sigma_{2}^{-1}(X-\mu)$ is the sum of Chi-square random variables, i.e., $(X-\mu)^{\top}\Sigma_{2}^{-1}(X-\mu)\sim\sum_{i}^{n}\lambda_{i}\chi_{1_{i}}^{2}$, where $\{\chi^{2}_{1_{i}}\}_{i=1}^{n}$ are independent Chi-square distributions with one degree of freedom and $\lambda_{i}$’s are the generalized eigenvalues of the matrix pencil $(\Sigma_{1},\Sigma_{2})$. Consequently, the mean and the variance of $Y$ are $\sum_{i=1}^{n}\lambda_{i}$ and $2\sum_{i=1}^{n}\lambda_{i}$, respectively.

### v2

@@ 4.1 Sample allocations and convergence rates
To investigate the effect of sample allocation, we set the ratio between the evaluation samples and training samples to be $L^{p}$, with $p=\frac{1}{8},\frac{1}{3},\frac{3}{4}$ respectively, and each experiment is repeated 100 times. In the first study, we assume the model is known and the likelihood function can be evaluated exactly, and we estimate only the marginal distribution of $Y$. That is, we use $\widehat{\mathrm{EIG}}_{\mathrm{m}}$. We also compare the performance of transport map-based estimators with that of the NMC estimator, where the NMC estimator is computed using the optimal sample allocation discussed in [24]. We then plot the estimated EIG along with their bias, variance, and mean squared error (MSE) for the different estimators in Figure 1 and 2. We observe that when we set the ratio of $M$ and $N$ to be proportional to $L^{1/3}$, the MSE decreases most rapidly as the total number of samples increases. The empirical convergence rate also aligns with the theory presented in the previous section. Next, we show in Figures 3 and  4 that by fixing the prior and using transport maps to approximate the posterior density, i.e., setting the EIG estimator to $\widehat{\mathrm{EIG}}_{\mathrm{pos}}$, we achieve the same rate.
@@
We then study the likelihood-free case, where we estimate either the likelihood and the marginal densities and obtain $\widehat{\mathrm{EIG}}_{\mathrm{lik}}$ (Figure 5), or the posterior and prior densities and obtain $\widehat{\mathrm{EIG}}_{\mathrm{pr}}$ (Figure 7). As shown in Figures 6 and 8, in both cases, the convergence rate of the MSE of the transport map-based EIG estimator matches the theoretical rate ($O(L^{-1})$) in the asymptotic regime, which is faster than that of the NMC ($O(L^{-2/3})$).
@@
We then present a few observations: The convergence rate of the MSE with sample allocation $\frac{M}{N}\sim O(L^{1/3})$ is of order $L^{-1}$ in the asymptotic region across all four cases. For $L^{p}$, when $p>\frac{1}{2}$, we observe that the convergence rate of the MSE is considerably slower than for $p\leq\frac{1}{2}$ in all four cases, which supports our theoretical results developed in Section 2.3. From the violin plot, we see that all the transport map-based results have smaller variance than NMC, and with allocation ratios of $L^{1/8}$ and $L^{1/3}$, transport maps produce good results even with small sample sizes. Another noteworthy observation is that $\widehat{\mathrm{EIG}}_{\mathrm{lik}}$ and $\widehat{\mathrm{EIG}}_{\mathrm{pr}}$ exhibit the same values, up to numerical error (see Figures 5 and  7). This occurs because we are in the linear Gaussian case, and if we assume the mean has been shifted to 0, then the map can be computed from covariance matrix of the joint samples. In other words, we can obtain $\widehat{\pi}_{Y}$, $\widehat{\pi}_{Y|X}$, $\widehat{\pi}_{X}$, and $\widehat{\pi}_{X|Y}$ directly from the sample covariance matrix of the joint distribution. Therefore, we have exactly $\frac{\widehat{\pi}_{X|Y}(x|y)}{\widehat{\pi}_{X}(x)}=\frac{\widehat{\pi}_{Y|X}(y|x)}{\widehat{\pi}_{Y}(y)}$ for any $x$ and $y$.
@@ 4.2 Transport-based estimators of focused and full EIG
We then study a nonlinear example, Mössbauer spectroscopy, to demonstrate the flexibility of transport maps. Mössbauer spectroscopy is a technique that leverages the Mössbauer effect, which refers to the nearly recoil-free emission and absorption of gamma radiation in solids. This technique is used to discover isomer shifts, quadrupole splitting, and magnetic Zeeman splitting. A detailed description of the model can be found in [24] and [44]. The parameters in this model are “center”, “width”, “height”, and “offset”. The absorption peak is modeled using a Lorentzian profile, as per the standard parameterization described in [29]. This allows the number of detector counts $y_{i}$ at a given velocity $d$ to be expressed by the following equation,
@@
All densities are estimated using $N$ pairs of samples. We set the “ground truth” to be the EIG estimator obtained using NMC. To obtain a consistent estimator of EIG using Monte Carlo for the focused case, we employed layered multiple importance sampling (LMIS) from [24] with $1.58\times 10^{6}$ samples (5623 for the outer loops and 281 for the inner loops). For the full case, we used the NMC algorithm with $10^{10}$ samples, where the allocation of samples between the inner and outer loops is determined using the optimal allocation proposed in [24]. In this case, we have $3.83\times 10^{6}$ samples for the inner loop and 2610 for the outer loop. We repeated the procedure 10 times to calculate the mean and variance of the EIG estimators obtained using LMIS and NMC. The mean of the Monte Carlo estimators in the focused and full cases is $1.54$ and $4.52$ respectively, with corresponding variances of $4.73\times 10^{-5}$ and $4.59\times 10^{-4}$. We then obtained EIG estimators using transport maps with an increasing total number of samples: 500, 5000, and 50000, and with the number of training and evaluation samples split according to the optimal sample allocation $M/N\sim O(L^{1/3})$, as derived in the previous section. We repeated the experiments, and obtained $30$ transport map-based EIG estimators. We present the numerical results of both cases in Figures 10 and 11.
@@
Finally, we present the joint distribution after applying the transport maps in Figure 12(a). As seen, the joint distribution after the push-forward map is very close to a standard normal distribution. To further illustrate the Gaussianity of the push-forward samples, we present the quantile-quantile plot (Q–Q plot) of the push-forward samples against that of a standard normal. To be precise, we generate 20 random vectors of unit length, and project 7-dimensional data onto 20 random directions, respectively. We then plot the quantile of these random projections against that of a standard normal distribution. If the push-forward distribution is exactly Gaussian, the Q–Q plot of the random projections would align with the black line (see Box 2.4 in [57]). As shown in Figure 12(b), only the tails of the push-forward samples deviate slightly, demonstrating the effectiveness of transport maps in transforming highly non-Gaussian samples into a standard Gaussian distribution.
@@
Figure 11: Convergence of different EIG estimators in the full case: (a) $\widehat{\mathrm{EIG}}_{\mathrm{pr}}$; (b) $\widehat{\mathrm{EIG}}_{\mathrm{lik}}$; (c) $\widehat{\mathrm{EIG}}_{\mathrm{m}}$ and $\widehat{\mathrm{EIG}}_{\mathrm{pos}}$. The dashed line represents the “ground truth” obtained using NMC.
@@ 4.3 EIG estimation in high dimensions: linear elasticity inverse problem
In this problem, we compare the proposed dimension reduction method for EIG estimation, referred to as Conditional Mutual Information (CMI), to alternative methods based on Principal Component Analysis (PCA) [30] and Canonical Correlation Analysis (CCA) [31]. We consider the problem of estimating mutual information between the Young’s modulus field in a wrench-shaped domain and a noisy observation of its displacement on the upper boundary after applying an external force to a subset of the boundary $\partial\Omega$. Following the setup in [4, 39, 58], the displacement $u:\Omega\rightarrow\mathbb{R}^{2}$ is related to the Young’s modulus $E:\Omega\rightarrow\mathbb{R}_{\geq 0}$ through the partial differential equation
@@
$\Sigma_{X}u^{\operatorname{PCA}}$ $=\lambda_{X}u^{\operatorname{PCA}}$
$\Sigma_{Y}v^{\operatorname{PCA}}$ $=\lambda_{Y}v^{\operatorname{PCA}}.$
@@
We then transform the samples $x^{i}$ and $y^{i}$ using these eigenvectors $\{u^{\operatorname{CMI}},v^{\operatorname{CMI}}\},\{u^{\operatorname{PCA}},v^{\operatorname{PCA}}\}$, and $\{u^{\operatorname{CCA}},v^{\operatorname{CCA}}\}$, respectively, where $\{u^{\operatorname{CMI}},v^{\operatorname{CMI}}\}$ is computed following Algorithm 2. The maps are trained using $N=2500$ samples. We then compute the Monte Carlo average using $M=10000$ evaluation samples to obtain the estimator $\frac{1}{M}\sum_{i=1}^{M}\frac{\widehat{\pi}_{X_{r}|Y_{s}}(x_{r}^{i}|y_{s}^{i})}{\pi_{X_{r}}(x_{r}^{i})}$. To illustrate the effectiveness of our proposed nonlinear dimension reduction method, we compute EIG as if the data were Gaussian. To be precise, we compute the EIG using the formula $\frac{1}{2}\log\frac{\det(\widehat{\Sigma}_{X})\det(\widehat{\Sigma}_{Y})}{\det(\widehat{\Sigma}_{X,Y})}$. Here, $\widehat{\Sigma}_{X},\widehat{\Sigma}_{Y}$ and $\widehat{\Sigma}_{X,Y}$ denote the sample covariance of $X,Y$ and the joint distribution of $X$ and $Y$, respectively. In practice, covariance matrices are approximated using their empirical counterparts estimated using $10000$ samples. For reproducibility, the same experiment is repeated $10$ times, and we show the error bars, representing two standard error of the mean in Figure 14.
@@
Figure 14: EIG estimates using transport maps and Gaussian, with truncated dimensions using CMI, PCA, and CCA for $r=s=1$ to $8$
@@ 5 Conclusion
In this paper, we study a key problem in Bayesian optimal experimental design — EIG estimation. We introduced the transport-map based method for EIG estimation in Section 2. This enables us to estimate EIG with both explicit and implicit models. In Section 2.3, we study the optimal sample allocation between the training and the evaluation samples for the purpose of minimizing the MSE. This is further verified with the linear Gaussian example presented in Section 4.1. In Section 3, we adopt the optimal dimension reduction technique from [4], enabling EIG to be estimated in high-dimensional settings. The numerical example presented in  Section 4 show that our proposed method performs similarly comparing to nested Monte Carlo in terms of the EIG but with a much cheaper cost. We also demonstrate the performance of our method in the high-dimensional setting in Section 4.3. The dimension reduction technique proposed in this paper outperform PCA, CCA, and its Gaussian counterpart in terms of EIG estimation.
@@ Appendix B Proofs of the main results
This quantity can be estimated using samples, i.e., $\frac{1}{M}\sum_{i=1}^{M}\log\frac{\pi_{Y|X}(y^{i}|x^{i})}{\pi_{Y}(y^{i})}$.
@@
We then study the covariance term $\mathbb{C}ov\left(\frac{1}{M}\sum_{i=1}^{M}\log\pi_{Y|X}(y^{i}|x^{i}),\frac{1}{M}\sum_{i=1}^{M}\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right)$. Note that $\mathbb{C}ov\left(\log\pi_{Y|X}(y^{i}|x^{i}),\log q_{Y}(y^{j};\widehat{\alpha}_{N})\right)=0$ for $i\neq j$. Therefore the expression can be simplified to
$\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{C}ov\left(\log\pi_{Y|X}(y^{i}|x^{i}),\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right).$
Using Cauchy-Schwarz, we have
$\mathbb{C}ov\left(\log\pi_{Y|X}(y^{i}|x^{i}),\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right)$
$\leq$ $\sqrt{\mathbb{V}\left[\log\pi_{Y|X}(y^{i}|x^{i})\right]\mathbb{V}\left[\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right]}.$
From previous analysis, the first term $\mathbb{V}\left[\log\pi_{Y|X}(y^{i}|x^{i})\right]$ is a constant, and the second term $\mathbb{V}\left[\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right]$ converges at a rate of $O(1+\frac{1}{N^{2}})$ using the law of total variance. Therefore,
$\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{C}ov\left(\log\pi_{Y|X}(y^{i}|x^{i}),\log q_{Y}(y^{i};\widehat{\alpha}_{N})\right)\sim\frac{1}{M}O\left(1+\frac{1}{N^{2}}\right)=O\left(\frac{1}{M}\right).$
To conclude, we have
@@
This theorem can be proved using the same techniques as in proof, thus omitted.
Proof B.3 (Proof of Theorem 2.3).
@@
Proof B.4 (Proof of Corollary 2.4).
@@ Appendix C Chi-square distribution
Suppose $X\in\mathbb{R}^{n}$ is normally distributed, i.e., $X\sim\mathcal{N}(\mu,\Sigma_{1})$, then the random variable $Y=(X-\mu)^{\top}\Sigma_{2}^{-1}(X-\mu)$ is the sum of Chi-square random variables, i.e., $(X-\mu)^{\top}\Sigma_{2}^{-1}(X-\mu)\sim\sum_{i=1}^{n}\lambda_{i}\chi_{1_{i}}^{2}$, where $\{\chi^{2}_{1_{i}}\}_{i=1}^{n}$ are independent Chi-square distributions with one degree of freedom and $\lambda_{i}$’s are the generalized eigenvalues of the matrix pencil $(\Sigma_{1},\Sigma_{2})$. Consequently, the mean and the variance of $Y$ are $\sum_{i=1}^{n}\lambda_{i}$ and $2\sum_{i=1}^{n}\lambda_{i}$, respectively.